Sol Summary – May 3th

Crew 211 Sol 12 Summary Report 3-May-2019

Sol: 12

Summary Title: Last day on Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The commander and vice-commander left the Hab at 7:00 am for an engineering EVA to get back the cubelanders and the UWB antennas. We had then a nice breakfast altogether before the morning EVA to North Ridge. Part of the team started to pack their cases until noon. The GreenHab officer and the crew journalist finalized their experiment early in the afternoon while the rest of the team started to clean the different modules. Half of the crew left at 4:00 pm for the last EVA to North Ridge looking at a beautiful sunset over the MDRS campus.

Look Ahead Plan: We plan to leave at 7:00 am with the crew car to get part of the crew 212 at Grand Junction. We should get back to the MDRS around noon and have a lunch altogether. The afternoon will be devoted to the handover.

Anomalies in work: No issues.

Weather: Sunny, almost no wind

Crew Physical Status: Good

EVA: The two EVAs of today allowed us to further explore the area to look for potential places to set-up a temporary antennas system for geolocation and information transmission. Kissing Camel, North Ridge and the Hab Ridge are definitely interesting places for such a project.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Report Summary report, Daily Photo Report

Support Requested: None

Sol Summary- May 02nd

Crew 211 Sol 11 Summary Report 2-May-2019

Sol: 11

Summary Title: Sunset on Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The commander and vice-commander began the day with a nice morning EVA along the Hab Ridge Road. After the breakfast, the commander and vice-commander deployed their respective experiment during an engineering EVA while the rest of the team finalized their own experiments. The deployment was postponed to this morning as different minor bugs prevented us to do it yesterday evening.The GreenHab officer harvested the Spirulina that was growing since more than a week and started to dry it. Part of the crew left at 5 pm for an EVA to Grey Moon. Tonight we will all observe the stars, enjoying the clear sky that the Utah offers.

Look Ahead Plan: We planned two EVAs for tomorrow as it is our last day on Mars.

Part of the crew will go to North Ridge in the morning. The rest of the team will end their experiment and start writing the summary report. We will get back the antennas and the cubelanders, in the morning during an engineering EVA. In the afternoon, we will do a second EVA, again to North Ridge. The rest of the afternoon will be devoted to packing our cases.

Anomalies in work: No issues.

Weather: Sunny, almost no wind

Crew Physical Status: Good

EVA: The morning EVA along Hab Ridge Road was shortened by the presence of steam in our helmets due to the low temperature at this time of the day. We didn’t manage to find the sinkhole but we had nevertheless the occasion to take nice pictures with the morning light. The second EVA to Grey moon started at 5 pm. The EVA team saw breathtaking panoramas.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, EVA Report, Daily Photo Report

Support Requested: None

Sol Summary – May 1st

Crew 211 Sol 10 Summary Report 1-May-2019

Sol: 10

Summary Title: At the crossroads

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After a difficult night with the storm and a temperature drop of more than 10 °c, we made some nice pancakes to comfort us. We then all worked on our experiments as the morning EVA was postponed to the early afternoon to wait for the mud to dry. Half of the crew made an EVA to Kissing Camel from 1:00 pm to 3:00 pm. The rest of the crew left at 4:30 pm for another EVA to Condor Chasma. The objective of the first EVA was to gather information about potential places to set-up a temporary antennas system for geolocation and information transmission. This will be one of the projects for the next UCL to Mars crew for the season 2019-2020. We end the day by deploying the antenna of the vice-commander experiment and the cubelanders of the commander experiments.

Look Ahead Plan: We plan to do another EVA as the end of the stays is approaching. We will get back the antennas and the cubelanders, in the morning, before walking north following the path 1103 to the sinkhole near Hab ridge road. A second EVA is planned in the afternoon to Grey Moon. We will also try to finalize our respective experiments in order to be able to write the final report for Friday night.

Anomalies in work: No issues.

Weather: Mostly sunny, almost no wind

Crew Physical Status: Good

EVA: We had two EVAs today to enjoy the nice weather after two days of showers. The EVA to Kissing Camel allowed us to prospect for good places to setup antennas to possibly create communication and geolocation systems in order to improve the communication between the Hab and EVA teams.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, EVA Report, Daily Photo Report

Support Requested: None

Sol Summary – April 30th

Crew 211 Sol 9 Summary Report 30-APR-2019

Sol: 9

Summary Title: Storm over Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: As we had numerous showers during the night, the morning EVA was unfortunately postponed to the afternoon and then cancelled as the weather deteriorated in the afternoon. We therefore worked all day on our experiments. We had an interesting meeting after the lunch to discuss with the crew journalist some articles of our Martian constitution. The experiment of the Commander and the Vice-Commander are now almost ready for field experiment.

Look Ahead Plan: Depending on the weather, we will try again to do a morning EVA to Kissing Camel and an afternoon EVA to Candor Chasma, as the end of our stay is approaching. We will in parallel work on our respective experiments and hopefully be able to start in the afternoon the field experiments with the cubelanders and the geolocation antennas.

Anomalies in work: No issues.

Weather: Showers, mostly cloudy and very windy

Crew Physical Status: Good

EVA: The EVA has been cancelled due to bad weather.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, Daily Photo Report

Support Requested: None

Sol Summary – April 29th

Crew 211 Sol 8 Summary Report 29-APR-2019

Sol: 8

Summary Title: Mars on TV

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After a quick breakfast, the commander welcomed the PBS team in spacesuit outside of the Hab. We then made a tour of the MDRS campus with the TV crew, each member of the UCL to Mars team presenting part of the infrastructures and discussing his/her experiment. Half of the crew left then for an EVA with Liz Adeola to Cow Dung Road between the Robert’s Rock Garden and Zurbin’s head. The EVA team got back at noon and we all had a nice lunch together. The crew commander and the crew astronomer were then interviewed for about one hour. All the crew members went then back to work once the TV crew left the campus.

Look Ahead Plan: Depending on the weather, we plan to start the day with an EVA to Kissing Camel. We will then work on our respective experiments for the rest of the day preparing the first experiment related EVA for the commander and vice-commander which should be organized on Wednesday.

Anomalies in work: No issues.

Weather: Showers, mostly cloudy and windy

Crew Physical Status: Good

EVA: Half of the crew explored the west part of Cow Dung Road with the PBS crew. Liz Adeola was part of the EVA team, as we her a spacesuit for the occasion. The EVA lasted for about an hour. The PBS team did a few drone footages and interviewed the EVA team members.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Sol Summary – April 27th

Crew 211 Sol 6 Summary Report 27-APR-2019

Sol: 6

Summary Title: No sign of ancient Martian life

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After three attempts, we finally had a successful EVA to Special Region. The rovers worked nominally and we spent one hour in Special region to explore the surrounding and try to find traces of ancient Martian life. Unfortunately, we didn’t find any traces of microbial or more evolved life. After a nice lunch, we all returned to our experiments until the CapCom.

Look Ahead Plan: An EVA is planned in the morning to go to White Rock Canyon on the 1101 and then go back to Cow Dung Road between the Robert’s Rock Garden and Zubrin’s head. We have been there for the first EVA and members of the crew would like to go back to get more soil samples. Half of the crew will stay at the base to work on their experiments. The rest of the day will be devoted to these experiments.

Anomalies in work: No issues.

Weather: Sunny in the morning and partly cloudy and windy in the afternoon.

Crew Physical Status: Good

EVA: Half of the crew managed to get to the Special region for a very nice EVA. After a one hour walk, we didn’t find any trace of ancient Martian life. However, we found a cavity which could be the entrance of a larger underground structure which may provide shelter to the crew. The cavity was not large enough to consider the possibility of creating an underground facility.

Reports to be filed: Journalist Report, GreenHab Report, HSO report, Science Report, Operations Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Sol Summary – April 25th

Good evening Atila,

Please find below the Sol 4 Summary.

Crew 211 Sol 4 Summary Report 25-APR-2019

Sol: 4

Summary Title: Rescuing a Martian

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The day started with an EVA to explore the Special Region. Unfortunately the battery level of the Spirit rover decreased to fast so we couldn’t reach the Special region. After a short briefing with the Hab, we decided to explore instead the Galileo road 1104 for about 30 minutes. We then worked on our experiments until the lunch. After a nice lunch altogether, the HSO performed a quick check-up on the crew member who twisted his ankle three days ago. As the ankle was still pretty much swelled, we decided with Dr Rupert that it would be better to see a doctor. We therefore drove the crew car to the Bicknell Health Care center. Fortunately, the ankle was not broken and the crew member should be healed within a week. The rest of the team worked during the afternoon on their respective experiments.

Look Ahead Plan: An EVA is planned in the morning to explore Special Region. Half of the crew will stay at the base to work, as usual, on their experiments. The rest of the day will be devoted to these experiments.

Anomalies in work: No issues.

Weather: Hot, bright and sunny most of the day, but windy in the evening.

Crew Physical Status: Twisted ankle still bloated but should be healed within a week

EVA: The aim of today’s EVA was to explore the Special Region. Unfortunately the battery level of the Spirit rover decreased to fast so we couldn’t reach the Special region. After a short briefing with the Hab, we decided to explore instead the Galileo road 1104 for about 30 minutes.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Science Report, Operations Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Sol Summary – April 24th

Crew 211 Sol 3 Summary Report 24-APR-2019

Sol: 3

Summary Title: A day with Martians

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After the daily medical check-up and a nice breakfast with chocolate pancakes, members of the crew left at 9:00 am for an EVA to get back Opportunity still stuck near Candor Chasma. To avoid any other issues with the rover, the EVA team got directly back to the Hab and continued its EVA on foot to explore the hills surrounding the base. The surprise of today came at 10:15 am when we learned that the student visit scheduled for tomorrow would in fact start 15 minutes later. We spent the next four hours with two groups of students visiting the different infrastructures of the MDRS and presenting our research works. These four hours have been very enriching for both the students and the UCL to Mars team. After the lunch, we all worked on our experiments. Tonight will be dedicated to sky observations with the remote controlled telescope.

Look Ahead Plan: An EVA is planned in the morning to explore Special Region. Half of the crew will stay at the base to work, as usual, on their experiments. The rest of the day will be mainly devoted to these experiments.

Anomalies in work: No issues.

Weather: Hot, bright and sunny all day.

Crew Physical Status: Twisted ankle still bloated (See HSO reports)

EVA: The aim of this third EVA was to recover the Opportunity rover. Once recovered, we decided to explore on foot the hills surrounding the MDRS.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Science Report, Operations Report, Astronomy Report, EVA report, EVA Request, Daily Photo Report + Timelapse

Support Requested: None

Sol Summary – April 23rd

Crew 211 Sol 2 Summary Report 23-APR-2019

Sol: 2

Summary Title: Feeling like a true Martian

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After the daily check-up performed by the HSO as part of his experiment, we enjoy a nice breakfast together. An EVA was planned at 9 am but as the Curiosity rover was not fully charged, we decided to postpone it. Most of the crew members worked therefore on their respective experiments while the crew astronomer and crew scientist got their first shots of the sun. We then worked altogether during the noon break on the writing of a Martian constitution with the crew journalist. During the early afternoon, some crew members prepared the first successful Martian bread. Half of the crew left at 4 pm for the planned EVA to Candor Chasma to look for possible shelters in case of high solar activity.

Look Ahead Plan: An EVA is planned in the morning to get back Opportunity near Candor Chasma and take additional soil samples. Half of the crew will stay at the base to work on their experiments. The crew astronomer will continue to take pictures of the sun in the morning and schedule a few observations with the Celestron telescope to get images from quasars. This will allow her to infer their luminosity and determine the evolution of their activity. The rest of the day will be devoted to the experiments.

Anomalies in work: No issues.

Weather: Hot, bright and sunny all day.

Crew Physical Status: Good, yesterday’s twisted ankle is less painful.

EVA: The aim of this second EVA was to explore the Candor Chasma region and look for caves. Underground structures could indeed be very helpful to protect the crew from solar wind/radiation in the case of increasing solar activity and CMEs (Coronal Mass Ejections). Unfortunately no caves have been found and we lost Opportunity near Candor Chasma. We’ll try to get the rover back during tomorrow’s EVA.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Astronomer Report, Operations Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site